OpenStudy (anonymous):

What is the transformation that occurs to the equation y = 2x if the equation changes to y = -2x+1 − 1? a)The transformation moves the graph 1 unit to the left and 1 unit down. b)The transformation moves the graph 1 unit to the left, 1 unit down and reflects that graph across the y-axis. c)The transformation moves the graph 1 unit to the left, 1 unit down and reflects that graph across the x-axis. d)The transformation moves the graph 1 unit to the right, 1 unit down and reflects that graph across the x-axis.

OpenStudy (imer):

I am pretty sure you are missing parentheses/brackets on the equation, am I right?

OpenStudy (anonymous):

yeah sorry, is y=2^x and y=-2^x+1 -1

OpenStudy (anonymous):

y equals negative two to the exponents x plus 1, minus 1

OpenStudy (imer):

ah I got it you meant y=-2^(x+1)-1, the answer should be "C".

OpenStudy (anonymous):

yeah thanks :)